# A Fourier-Jacobi Dirichlet series attached to modular forms of $\textup{SO}(2,4)$

### v1

@@ 2. Preliminaries
respectively, and put $G^{+}(V):=G(V)\cap A^{+}(V)$, the even Clifford group. We also define the homomorphisms
$\displaystyle\begin{split}\tau:G^{+}(V)&\longrightarrow\textup{SO}^{\phi}(V)\\ \alpha&\longmapsto\tau(\alpha)x:=\alpha x\alpha^{-1}.\end{split}$ (2)
@@
$\displaystyle\begin{split}\nu:G^{+}(V)&\longrightarrow F^{\times}\\ \alpha&\longmapsto\alpha\alpha^{*}.\end{split}$ (3)
It is well known that $\tau$ is surjective with kernel $F^{\times}$.
Let now $\mathfrak{g}$ denote a maximal order in $F$, i.e. $\mathfrak{g}=\mathbb{Z}$ or $\mathbb{Z}_{p}$, depending on $F=\mathbb{Q}$ or $\mathbb{Q}_{p}$. By a $\mathfrak{g}$-lattice $L$ in $V$, we mean a finitely generated $\mathfrak{g}$-module that spans $V$ over $F$. It is called $\mathfrak{g}$-integral if $\phi[x]\in\mathfrak{g}$ for all $x\in L$. We say $L$ is $\mathfrak{g}$-maximal if it maximal among all integral $\mathfrak{g}$-lattices. We also define its dual lattice via
@@
The level $q$ of $L$ is defined as the least positive integer such that $q\phi[x]\in\mathfrak{g},\ \forall x\in L^{*}$. We also define the integral orthogonal group via
@@
$L[q,\mathfrak{b}]:=\{x\in V\mid\phi[x]=q\textup{ and }\phi(x,L)=\mathfrak{b}\}.$ (4)
@@ 3. Modular Forms on $\textup{SO}(2,n)$
of real signatures $(1,n+1)$ and $(2,n+2)$, respectively. Let also $V_{0}:=\mathbb{Q}^{n+2}$ and $V_{1}:=\mathbb{Q}^{n+4}$ and consider the quadratic spaces $(V_{0},\phi_{0}),\ (V_{1},\phi_{1})$, where
@@
for $i=0,1$. We then have that $\phi:=\phi_{0}\mid_{V\times V}$ is just $(x,y)\longmapsto-x^{t}Sy/2$, and we make the assumption that $L=\mathbb{Z}^{n}$ is a maximal $\mathbb{Z}$-lattice with respect to $\phi$.
From [Shi04, Lemma 6.3], we then obtain that $L_{0}:=\mathbb{Z}^{n+2}$ is a $\mathbb{Z}$-maximal lattice in $V_{0}$.
If now $K\supset\mathbb{Q}$ is a field, the corresponding orthogonal groups of $K$-rational points are given by
$\textup{SO}^{\phi_{0}}(V_{0}\otimes K):=\{g\in\textup{SL}_{n+2}(K)\mid g^{t}S_{0}g=S_{0}\},$
$\textup{SO}^{\phi_{1}}(V_{1}\otimes K):=\{g\in\textup{SL}_{n+4}(K)\mid g^{t}S_{1}g=S_{1}\}.$
Let now $\mathcal{H}_{S}$ denote one of the connected components of
@@
$\mathcal{H}_{S}:=\{z=x+iy\in V_{0}\otimes_{\mathbb{R}}\mathbb{C}\mid y\in\mathcal{P}_{S}\}.$ (5)
@@
with $A\in M_{n+2}(\mathbb{R}),\alpha,\beta,\gamma,\delta\in\mathbb{R}$ and $a,b,c,d$ real column vectors. Now the map
$Z\longmapsto g\langle Z\rangle=\frac{-\frac{1}{2}S_{0}[Z]b+AZ+c}{-\frac{1}{2}S_{0}[Z]\gamma+d^{t}Z+\delta}$ (6)
gives a well-defined transitive action of $\textup{SO}^{\phi_{1}}(V_{1}\otimes\mathbb{R})^{+}$ on $\mathcal{H_{S}}$, where $\textup{SO}^{\phi_{1}}(V_{1}\otimes\mathbb{R})^{+}$ denotes the connected component of the identity of $\textup{SO}^{\phi_{1}}(V_{1}\otimes\mathbb{R})$. The denominator of the above expression is the factor of automorphy
@@
Let now $L_{1}:=\mathbb{Z}^{n+4}$ and define the groups
$\Gamma(L_{0}):=\{g\in\textup{SO}^{\phi_{0}}(V_{0})\mid gL_{0}=L_{0}\},$
$\Gamma(L_{1}):=\{g\in\textup{SO}^{\phi_{1}}(V_{1})\mid gL_{1}=L_{1}\}.$
Let also $\Gamma^{+}(L_{0})=\Gamma(L_{0})\cap\textup{SO}^{\phi_{0}}(V_{0}\otimes\mathbb{R})^{+}$. Moreover, we set
$\Gamma_{S}:=G_{\mathbb{R}}^{0}\cap\Gamma(L_{1}),$
and
$\widehat{\Gamma}_{S}:=\{M\in\Gamma_{S}\mid M\in 1_{n+4}+\textup{M}_{n+4}(\mathbb{Z})S_{1}\},$
the discriminant kernel.
@@
Let $k\in\mathbb{Z}$ and $\widehat{\Gamma}_{S}\leq\Gamma\leq\Gamma_{S}$ a subgroup of finite index. A holomorphic function $F:\mathcal{H}_{S}\longrightarrow\mathbb{C}$ is called a modular form of weight $k$ with respect to $\Gamma$ if it satisfies the equation

### v2

@@ 2. Preliminaries
respectively, and put $G^{+}(V):=G(V)\cap A^{+}(V)$, the even Clifford group. We define the homomorphisms
(2.1) $\displaystyle\begin{split}\tau:G^{+}(V)&\longrightarrow\textup{SO}^{\phi}(V)\\ \alpha&\longmapsto\tau(\alpha)x:=\alpha x\alpha^{-1}.\end{split}$
@@
(2.2) $\displaystyle\begin{split}\nu:G^{+}(V)&\longrightarrow F^{\times}\\ \alpha&\longmapsto\alpha\alpha^{*}.\end{split}$
It is well known that $\tau$ is surjective with kernel $F^{\times}$. We also define the spinor norm
$\displaystyle\sigma:\textup{SO}^{\phi}(V)\longrightarrow\mathbb{Q}^{\times}/(\mathbb{Q}^{\times})^{2}$
as follows. If $\alpha\in\textup{SO}^{\phi}(V)$, we pick an element $\xi\in G^{+}(V)$ such that $\tau(\xi)=\alpha$ and denote by $\sigma(\alpha)$ the coset $\nu(\xi)(\mathbb{Q}^{\times})^{2}$, which is determined by $\alpha$.
Let now $\mathfrak{g}$ denote a maximal order in $F$, i.e. $\mathfrak{g}=\mathbb{Z}$ or $\mathbb{Z}_{p}$, depending on $F=\mathbb{Q}$ or $\mathbb{Q}_{p}$. By a $\mathfrak{g}$-lattice $L$ in $V$, we mean a finitely generated $\mathfrak{g}$-module that spans $V$ over $F$. It is called $\mathfrak{g}$-integral if $\phi[x]\in\mathfrak{g}$ for all $x\in L$. We say $L$ is $\mathfrak{g}$-maximal if it is maximal among all integral $\mathfrak{g}$-lattices. We also define its dual lattice via
@@
The level $q$ of $L$ is defined as the least positive integer such that $q\phi[x]\in\mathfrak{g},\ \forall x\in L^{*}$. A vector $v\in L$ is called primitive if
$\forall k\in\mathbb{Z},w\in L:(v=kw\implies k=\pm 1).$
We also define the integral orthogonal group via
@@
(2.3) $L[q,\mathfrak{b}]:=\{x\in V\mid\phi[x]=q\textup{ and }\phi(x,L)=\mathfrak{b}\}.$
@@ 3. Modular Forms on $\textup{SO}(2,n)$
of real signatures $(1,n+1)$ and $(2,n+2)$, respectively.
Let also $V_{0}:=\mathbb{Q}^{n+2}$ and $V_{1}:=\mathbb{Q}^{n+4}$ and consider the quadratic spaces $(V_{0},\phi_{0}),\ (V_{1},\phi_{1})$, where
@@
for $i=0,1$.
We then have that $\phi:=\phi_{0}\mid_{V\times V}$ is just $(x,y)\longmapsto-x^{t}Sy/2$, and we make the assumption that $L=\mathbb{Z}^{n}$ is a maximal $\mathbb{Z}$-lattice with respect to $\phi$.
From [shimura2004arithmetic, Lemma 6.3], we then obtain that $L_{0}:=\mathbb{Z}^{n+2}$ is a $\mathbb{Z}$-maximal lattice in $V_{0}$.
If now $R$ is a $\mathbb{Q}$-algebra, the corresponding orthogonal groups of $R$-rational points are given by
$G_{R}^{*}=\textup{SO}^{\phi_{0}}(V_{0}\otimes R):=\{g\in\textup{SL}_{n+2}(R)\mid g^{t}S_{0}g=S_{0}\},$
$G_{R}=\textup{SO}^{\phi_{1}}(V_{1}\otimes R):=\{g\in\textup{SL}_{n+4}(R)\mid g^{t}S_{1}g=S_{1}\}.$
In particular, for each prime $p$, we will abbreviate by $G_{p}$ the group $G_{\mathbb{Q}_{p}}$. We will write $G_{\mathbb{A}}$ for the adelisation of $G$, $G_{\mathbb{R}}$ for the infinite part of $G_{\mathbb{A}}$ and $G_{\mathbb{A},f}$ for its finite part. We use analogous notation for $G^{*}$.
We now collect some standard facts about automorphic forms on $G$ (see for example [sugano] or [krieg_jacobi]). We let $\mathcal{H}_{S}$ denote one of the connected components of
@@
(3.1) $\mathcal{H}_{S}:=\{z=x+iy\in V_{0}\otimes_{\mathbb{R}}\mathbb{C}\mid y\in\mathcal{P}_{S}\}.$
@@
with $A\in\textup{M}_{n+2}(\mathbb{R}),\alpha,\beta,\gamma,\delta\in\mathbb{R}$ and $a,b,c,d$ real column vectors. Now the map
(3.2) $Z\longmapsto g\langle Z\rangle=\frac{-\frac{1}{2}S_{0}[Z]b+AZ+c}{-\frac{1}{2}S_{0}[Z]\gamma+d^{t}Z+\delta}$
gives a well-defined transitive action of $G_{\mathbb{R}}^{+}$ on $\mathcal{H_{S}}$, where $G_{\mathbb{R}}^{+}$ denotes the connected component of the identity of $G_{\mathbb{R}}$. The denominator of the above expression is the factor of automorphy
@@
Let $Z_{0}$ denote any point of $\mathcal{H}_{S}$ with real part $0$ and denote by $D_{\mathbb{R}}$ its stabiliser in $G_{\mathbb{R}}^{+}$. Then, we have that $G_{\mathbb{R}}^{+}/D_{\mathbb{R}}\cong\mathcal{H}_{S}$.
In the following, we will need the notion of an automorphic form. For each prime number $p<\infty$, let $D_{p}:=G_{p}\cap\textup{SL}_{n+4}(\mathbb{Z}_{p})$ and let
$D_{f}:=\prod_{p<\infty}D_{p}.$
@@
Let $k\geq 0$. A function $\bm{F}:G_{\mathbb{A}}\longrightarrow\mathbb{C}$ is called a holomorphic cusp form of weight $k$ with respect to $D_{f}$ if the following conditions hold:
  1. (1)
    $\bm{F}(\gamma gu)=\bm{F}(g)\textup{ }\forall\gamma\in G_{\mathbb{Q}},u\in D_{f}$.
  2. (2)
    For any $g=g_{\infty}g_{f}$ with $g_{\infty}\in G_{\mathbb{R}}^{+}$ and $g_{f}\in G_{\mathbb{A},f}$, $\bm{F}(g_{\infty}g_{f})j(g_{\infty},Z_{0})^{k}$ depends only on $g_{f}$ and $Z=g_{\infty}\langle Z_{0}\rangle$ and is holomorphic on $\mathcal{H}_{S}$ as a function of $Z$.
  3. (3)
    $\bm{F}$ is bounded on $G_{\mathbb{A}}$.
Denote the above space by $\mathfrak{S}_{k}(D_{f})$.
Let now $\Gamma$ denote a congruence subgroup of $G_{\mathbb{Q}}\cap G_{\mathbb{R}}^{+}$, in the sense of [Shimura_Euler_Product, p. 62]. We then have the following definition of an orthogonal modular form in the classical sense.
Definition 3.2.
A holomorphic function $F:\mathcal{H}_{S}\longrightarrow\mathbb{C}$ is called a modular form of weight $k\geq 0$ with respect to $\Gamma$ if it satisfies the equation